# MonoForce: Learnable Image-conditioned Physics Engine

### v1

@@ Abstract
We propose a novel model for the prediction of robot trajectories on rough offroad terrain from the onboard camera images. This model enforces the laws of classical mechanics through a physics-aware neural symbolic layer while preserving the ability to learn from large-scale data as it is end-to-end differentiable. The proposed hybrid model integrates a black-box component that predicts robot-terrain interaction forces with a neural-symbolic layer. This layer includes a differentiable physics engine that computes the robot’s trajectory by querying these forces at the points of contact with the terrain. As the proposed architecture comprises substantial geometrical and physics priors, the resulting model can also be seen as a learnable physics engine conditioned on real images that delivers $10^{4}$ trajectories per second. We argue and empirically demonstrate that this architecture reduces the sim-to-real gap and mitigates out-of-distribution sensitivity. The differentiability, in conjunction with the rapid simulation speed, makes the model well-suited for various applications including model predictive control, trajectory shooting, supervised and reinforcement learning or SLAM. The codes and data are publicly available [^1] https://github.com/ctu-vras/monoforce.
@@ I Introduction
Autonomous robotics in off-road environments holds immense promise for a variety of commercial applications, including outdoor logistics, inspections, and forestry operations. Yet, unlike systems designed for controlled environments, such as factories, off-road autonomous systems still remain prohibitively immature for dependable deployment. The main challenge that prevents reliable off-road deployment lies in the inability to accurately predict the behavior of the robot on the terrain. For example, the prediction of robot trajectory in tall vegetation or muddy and rocky terrains, especially from camera images and lidar scans, remains an unresolved problem in outdoor robotics.
Figure 1: Qualitative results: The first three rows show (i) onboard camera images (the only input), (ii) predicted supporting terrain and (iii) 3D view with predicted trajectory and visualized ground truth lidar point clouds (not used for estimation). While the majority of rigid objects are correctly reconstructed from the camera image, robot flippers, tall grass, fallen branches or soft hanging branches are correctly suppressed as they have no influence on the resulting robot trajectory. Predicted terrain friction, dampening, and stiffness are not visualized for bravity.
@@
Over the last decade, roboticists proposed a wide variety of white-box [1, 2, 3] and black-box models [4, 5, 6, 7, 8] for the offroad trajectory prediction. Black-box models primarily suffer from a severe out-of-distribution problem – the phenomenon where the distribution of training data does not correspond to the testing data. This problem naturally arises from the fact that the training data includes only the trajectories from safe terrain traversals, such as small terrain steps, while the decision about terrain traversability also naturally comprises the non-traversable terrains, such as tall cliffs, Figure 4. Since the step-to-cliff-generalization of large state-of-the-art black-box models is typically poor, the safety of the resulting system is arbitrarily inferior. On the other hand, white-box architectures offer good generalization due to substantial inductive bias but often suffer from the sim-to-real gap – the phenomenon where a model """"trained in"""" or """"replaced by"""" simulation faces challenges or discrepancies when applied in the real world. Although several techniques, such as rapid motor adaptation [9], are known to partially suppress this issue through domain randomization, the sim-to-real gap, as well as the out-of-distribution problem, makes trajectory prediction from camera images prohibitively unreliable. We introduce a grey-box model that combines the best of both worlds to achieve better generalization and smaller sim-to-real gap. This model enforces white-box laws of classical mechanics through a physics-aware neural symbolic layer, while preserving the ability to learn from huge data as it is end-to-end differentiable. We demonstrate that even just a single onboard camera is enough for reliable autonomous deployment in off-road environments; see Figures 13 for some qualitative results and Figure 5 for quantitative evaluation.
@@
Figure 3: Qualitative results: Given input onboard image from robot’s camera (first row), the method distinguishes stones and trees as obstacle, while vegetation is suppressed (see 3D view in third row) the stone is preserved despite having almost the same green color as vegetation due to the moss coverage. Heightmap color encodes predicted terrain friction, dark blue arrows shows predicted contact forces. Predicted robot trajectory is visualized in light green. The ground truth robot motion is captured in the second row.
@@
We highlight that all state-of-the-art models, including the proposed one, are inherently statistically inconsistent as it is impossible to provide training data covering dangerous cases. Consequently, providing a close-to-infinite number of training examples, as typical, for large foundation models [10] does not imply a superior performance. Addressing the statistical inconsistency for the trajectory prediction task without resorting to real-world robot-damaging trajectories remains an open challenge. In this work, we make a small step toward this direction by introducing a grey-box model that leverages: (i) the strengths of black-box models to learn effectively from real data, thereby reducing the sim-to-real gap, and (ii) the strengths of white-box approaches to generalize effectively to unseen data, thereby addressing the out-of-distribution problem.
@@
As the proposed architecture comprises substantial geometrical and physics priors, the resulting model can be also seen as learnable physics engine conditioned by a real image that delivers $10^{4}$ trajectories per second; see Figure 2 for details (the engine efficiency is also investigated in the Section IV-D). In addition to that, the model is end-to-end-differentiable; therefore, gradients can be backpropagated towards its (i) convolutional filters, (ii) camera and robot parameters, (iii) control, (iv) positions and (v) terrain properties. The differentiability, in conjunction with the rapid simulation speed, draws the model suitable for a myriad of tasks, including model predictive control [11], trajectory shooting [12], supervised and reinforcement learning [13], SLAM [14], online robot-model reidentification or camera recalibration [15]. The explainable structure of the proposed architecture also delivers a variety of intermediate outputs, such as terrain shape and its physical properties, robot-terrain reaction forces or contact points, which can all serve as efficient sources of self-supervision if measured during the training set creation or restricted in PINN-like manner [16]. We observed that the instability of gradient computation in existing differentiable simulators, such as NVIDIA’s WARP [17] and Google’s BRAX [18], makes them prohibitively unreliable and slow for both the learning and the inference. As a suitable replacement, we introduce a differentiable neural symbolic layer, which involves our from-scratch implemented physics engine, into the proposed architecture. This solution outperforms existing works in gradient stability and computational speed while preserving sufficient accuracy for reliable trajectory prediction. As all target applications, including learning, control, planning, and SLAM, can be naturally parallelized, we also achieve significant speed-up through massive parallelization on GPU.
@@ II-A Black-box (Data-driven) Models
The environment representation modeling from a single [25] or multiple RGB images [22] in top-down view has been widely studied in the literature. The geometry of the scene prediction (taking into account occlusions) from stereo or depth-camera input has been addressed in [26]. The importance and benefits of sensor fusion (camera images, LiDAR and radar) for environment representation have been demonstrated in [27]. The authors of RoadRunner [28] developed an end-to-end learning-based framework that predicts the traversability and elevation maps from multiple images and a LiDAR voxel map. However, the robot-terrain interaction and trajectory prediction tasks are not considered in these works.
@@
Figure 7: Detailed view of MonoForce’s heightmaps: Geometrical heightmap replicates lidar heightmap. Soft heightmap outlines soft parts of the terrain (those which does not generate any robot-terrain interaction forces) such as vegetation. Terrain heightmap consists of two layers: (i) Supporting terrain: geometrical heightmap with subtracted soft parts of the terrain; (ii) Friction heightmap: see significantly higher values on dirt.
@@ II-B White-box (Physics-based) Models
The vehicle-terrain interaction modeling and traction mechanics has been widely studied in the literature [35][36]. To provide accurate analysis and prediction of off-road vehicles behavior on the terrain, a set of high-fidelity physics-based software tools have been developed. For example, Vortex Studio [37] allows for realistic wheeled and tracked locomotion simulations and extensive land and even planetary environments modeling. The Chrono [38] is a physics-based engine that is designed for ground vehicle-tyre-terrain interactions simulation. It supports the integration of various vehicle and tire models and can simulate deformable terrain. The AGX physics engine (more specifically, its terrain dynamics model agxTerrain[39] enables simulation of soil dynamics and interactions with heavy vehicles. This software delivers realistic, high-fidelity simulations for applications involving complex soil-tool interactions. However, the mentioned physics engines are not differentiable, which makes them unsuitable for end-to-end learning and integration of exteroceptive real sensor data.
@@ III Method
A detailed overview of the proposed architecture that converts images and control commands into trajectories is depicted in Figure 6. The model consists of several learnable modules that deeply interact with each other. The terrain encoder carefully transforms visual features from input image into the heightmap space using known camera geometry. The resultant heightmap features are further refined into interpretable physical quantities that capture properties of the terrain such as its shape, friction, stiffness and damping. Next, the physics engine combines the terrain properties with the robot model, robot state and control commands and delivers reaction forces at points of robot-terrain contacts. It then solves the equations of motion dynamics by integrating these forces and delivers the trajectory of the robot. Since the complete computational graph of the feedforward pass is retained, the backpropagation from an arbitrary loss, constructed on top of delivered trajectories, or any other intermediate outputs, is at hand.
@@ III-A Terrain Encoder
The part of the MonoForce architecture (Figure 6) that predicts terrain properties $\mathbf{m}$ from sensor measurements $\mathbf{z}$ is called terrain encoder. The proposed architecture starts by converting pixels from a 2D image plane into a heightmap with visual features. Since the camera is calibrated, there is a substantial geometrical prior that connects heightmap cells with the pixels. We incorporate the geometry through the Lift-Splat-Shoot architecture [22]. This architecture uses known camera intrinsic parameters to estimate rays corresponding to particular pixels – pixel-rays, Figure 13. For each pixel-ray, the convolutional network then predicts depth probabilities and visual features. Visual features are vertically projected on a virtual heightmap for all possible depths along the corresponding ray. The depth-weighted sum of visual features over each heightmap cell is computed, and the resulting multichannel array is further refined by deep convolutional network to estimate the terrain properties $\mathbf{m}$.
@@ III-D End-to-end Learning
Terrain loss that minimizes the difference between ground truth $\mathcal{H}_{t}^{\star}$ and predicted $\mathcal{H}_{t}$ supproting heightmaps containing rigid objects detected with Microsoft’s image segmentation model SEEM [34], that is derived from Segment Anything foundation model [46]:
@@
Since the architecture Figure 2 is end-to-end differentiable, we can directly learn to predict all intermediate outputs just using trajectory loss (8). An example of terrain learning with the trajectory loss is visualized in Figure 9. To make the training more efficient and the learned model to be explainable, we employ the geometrical loss (9) and terrain loss (10) as regularization terms. stat
@@
Figure 11: MonoForce prediction examples. Left: The robot is moving through a narrow passage between a wall and tree logs. Right: The robot is moving on a gravel road with rocks on the sides. It starts its motion from the position marked with a coordinate frame and the trajectory is predicted for $10~{}[$\sec$]$ using real control commands. The camera images are taken from the robot’s initial position (top row). The visualization includes predicted supporting terrain $\mathcal{H}_{t}$ (second row). It is additionally shown in 3D and colored with predicted friction values (third row). The estimated and ground-truth robot’s trajectories are visualized in the green and red colors, respectively. the two top rows show the predictions projected to the robot’s camera images
@@
The Figure 11 show the prediction examples of the MonoForce model in diverse outdoor environments. From the example on the left, we can see that the model correctly predicts the robot’s trajectory and the terrain shape suppressing traversable vegetation, while the rigid obstacles (wall and tree logs) are correctly detected. The example on the right demonstrates the model’s ability to predict the robot’s trajectory ($10~{}[$\sec$]$-long horizon) with reasonable accuracy and to detect the rigid obstacles (stones) on the terrain. It could also be noticed that the surfaces that provide the robot good traction (paved and gravel roads) are marked with a higher friction value, while for the objects that might not give a good contact with the robot’s tracks (walls and tree logs) the friction value is lower.
@@
We argue that the friction estimates are approximate and an interesting research direction could be to compare them with the real-world measurements or with the values provided from a high-fidelity physics engine (e.g. AGX Dynamics [39]). However, one of the benefits of our differentiable approach is that the model does not require ground-truth friction values for training. The predicted heightmap’s size is $12.8\times 12.8$\mathrm{m}$^{2}$ and the grid resolution is $0.1$\mathrm{m}$$. It has an upper bound of $1~{}[$\mathrm{m}$]$ and a lower bound of $-1~{}[$\mathrm{m}$]$. This constraint was introduced based on the robot’s size and taking into account hanging objects (tree branches) that should not be considered as obstacles (Figure 17). Additionally, the terrain is predicted in the gravity-aligned frame. That is made possible thanks to the inclusion of camera intrinsics and extrinsics as input to the model, Figure 6. It also allows correctly modeling the robot-terrain interaction forces (and thus model robot’s trajectory accurately) for the scenarios with non-flat terrain, for example, going uphill or downhill. This will not be possible if only camera images are used as input.
@@ IV-A Dataset: ROUGH
Figure 12: Two robot platforms used to collect the ROUGH dataset. Notice that one platform has only for massive flippers, the other has also two main tracks.
@@
Despite the amount and annotation quality of the data provided in the RELLIS-3D sequences, it lacks examples of a robot moving over hills, obstacles, and traversing high grass. To fill this gap in autonomous off-road navigation, we release a new Rugged Obstacles and Uneven Ground Hardcore (ROUGH) dataset, containing sensory data from forest and field scenarios recorded with the two robot platforms: MARV, Bluebotics AbsolemFigure 12. It contains several hours of driving with the mid-sized robots in challenging terrain. The ultimate goal is to provide recordings of traversals on (or through) flexible natural obstacles like tall grass, hanging tree branches, mud, dense undergrowth, etc. The ROUGH sequences are collected with shape-changing tracked robots (Figure 12), which allows capturing a much larger range of dynamic responses - simply moving the auxiliary tracks changes the center and moments of inertia and the contact surface. To correctly utilize the effects of dynamics, ROUGH provides not only the sensory data, but also robot models with dynamics properties. The data set also contains carefully time-synchronized point cloud scans from Ouster OS0-128 lidar and corresponding RGB images from $4$ Basler cameras installed on the robots (front, rear, left, and right). We also record robot localization data obtained using ICP SLAM [48]. It was a better choice than GPS which is unreliable near or under tree canopy and close to buildings. The controls ($\mathbf{u}$ in Figure 6) in the form of commanded tracks velocities and angles are recorded and used to train and evaluate the MonoForce model.
After training the MonoForce terrain encoder on the RELLIS-3D dataset, we fine-tune it on our ROUGH data sequences. Please note that the dataset is under construction at the moment of writing and will be extended with the help of new robot platforms and data collection scenarios. The ROUGH dataset will be made publicly available [^2] https://github.com/ctu-vras/rough-dataset.

### v2

@@ Abstract
We propose a novel model for the prediction of robot trajectories on rough offroad terrain from the onboard camera images. This model enforces the laws of classical mechanics through a physics-aware neural symbolic layer while preserving the ability to learn from large-scale data as it is end-to-end differentiable. The proposed hybrid model integrates a black-box component that predicts robot-terrain interaction forces with a neural-symbolic layer. This layer includes a differentiable physics engine that computes the robot’s trajectory by querying these forces at the points of contact with the terrain. As the proposed architecture comprises substantial geometrical and physics priors, the resulting model can also be seen as a learnable physics engine conditioned on real images that delivers $10^{4}$ trajectories per second. We argue and empirically demonstrate that this architecture reduces the sim-to-real gap and mitigates out-of-distribution sensitivity. The differentiability, in conjunction with the rapid simulation speed, makes the model well-suited for various applications including model predictive control, trajectory shooting, supervised and reinforcement learning, or SLAM. The codes and data are publicly available [^1] https://github.com/ctu-vras/monoforce.
@@ I Introduction
Autonomous robotics in off-road environments holds immense promise for various commercial applications, including outdoor logistics, inspections, and forestry operations. Yet, unlike systems designed for controlled environments, such as factories, off-road autonomous systems still remain prohibitively immature for dependable deployment. The main challenge that prevents reliable off-road deployment lies in the inability to predict the behavior accurately of the robot on the terrain. For example, the prediction of robot trajectory in tall vegetation or muddy and rocky terrains, especially from camera images and lidar scans, remains an unresolved problem in outdoor robotics.
Figure 1: Qualitative results: The first three rows show (i) onboard camera images (the only input), (ii) predicted supporting terrain and (iii) a 3D view with predicted trajectory and visualized ground truth lidar point clouds (not used for estimation). While the majority of rigid objects are correctly reconstructed from the camera image, robot flippers, tall grass, fallen branches, or soft hanging branches are correctly suppressed as they do not influence the resulting robot trajectory. Predicted terrain friction, dampening, and stiffness are not visualized for brevity.
@@
Over the last decade, roboticists proposed a wide variety of white-box [1, 2, 3] and black-box models [4, 5, 6, 7, 8] for the off-road trajectory prediction. Black-box models primarily suffer from a severe out-of-distribution problem – the phenomenon where the distribution of training data does not correspond to the testing data. This problem naturally arises from the fact that the training data includes only the trajectories from safe terrain traversals, such as small terrain steps, while the decision about terrain traversability also naturally comprises the non-traversable terrains, such as tall cliffs, Figure 4. Since the step-to-cliff-generalization of large state-of-the-art black-box models is typically poor, the safety of the resulting system is arbitrarily inferior. On the other hand, white-box architectures offer good generalization due to substantial inductive bias but often suffer from the sim-to-real gap – the phenomenon where a model """"trained in"""" or """"replaced by"""" simulation faces challenges or discrepancies when applied in the real world. Although several techniques, such as rapid motor adaptation [9], are known to partially suppress this issue through domain randomization, the sim-to-real gap, as well as the out-of-distribution problem, makes trajectory prediction from camera images prohibitively unreliable. We introduce a grey-box model that combines the best of both worlds to achieve better generalization and a smaller sim-to-real gap. This model enforces white-box laws of classical mechanics through a physics-aware neural symbolic layer, while preserving the ability to learn from huge data as it is end-to-end differentiable. We demonstrate that even just a single onboard camera is enough for reliable autonomous deployment in off-road environments; see Figures 13 for some qualitative results and Figure 5 for quantitative evaluation.
@@
Figure 3: Qualitative results: Given input onboard image from the robot’s camera (first row), the method distinguishes stones and trees as obstacles, while vegetation is suppressed (see 3D view in third row) the stone is preserved despite having almost the same green color as vegetation due to the moss coverage. Heightmap color encodes predicted terrain friction, and dark blue arrows show predicted contact forces. The predicted robot trajectory is visualized in light green. The ground truth robot motion is captured in the second row.
@@
We highlight that all state-of-the-art models, including the proposed one, are inherently statistically inconsistent as it is impossible to provide training data covering dangerous cases. Consequently, providing a close-to-infinite number of training examples, as typical, for large foundation models [10] does not imply a superior performance. Addressing the statistical inconsistency for the trajectory prediction task without resorting to real-world robot-damaging trajectories remains an open challenge. In this work, we make a step toward this direction by introducing a grey-box model that leverages: (i) the strengths of black-box models to learn effectively from real data, thereby reducing the sim-to-real gap, and (ii) the strengths of white-box approaches to generalize effectively to unseen data, thereby addressing the out-of-distribution problem.
@@
As the proposed architecture comprises substantial geometrical and physics priors, the resulting model can be also seen as learnable physics engine conditioned by a real image that delivers $10^{4}$ trajectories per second; see Figure 2 for details (the engine efficiency is also investigated in the Section IV-D). In addition to that, the model is end-to-end-differentiable; therefore, gradients can be backpropagated towards its (i) convolutional filters, (ii) camera and robot parameters, (iii) control, (iv) positions, and (v) terrain properties. The differentiability, in conjunction with the rapid simulation speed, makes the model suitable for a myriad of tasks, including model predictive control [11], trajectory shooting [12], supervised and reinforcement learning [13], SLAM [14], online robot-model reidentification or camera recalibration [15]. The explainable structure of the proposed architecture also delivers a variety of intermediate outputs, such as terrain shape and its physical properties, robot-terrain reaction forces or contact points, which can all serve as efficient sources of self-supervision if measured during the training set creation or restricted in PINN-like manner [16]. We observed that the instability of gradient computation in existing differentiable simulators, such as NVIDIA’s WARP [17] and Google’s BRAX [18], makes them prohibitively unreliable and slow for both the learning and the inference. As a suitable replacement, we introduce a differentiable neural symbolic layer, which involves our from-scratch implemented physics engine, into the proposed architecture. This solution outperforms existing works in gradient stability and computational speed while preserving sufficient accuracy for reliable trajectory prediction. As all target applications, including learning, control, planning, and SLAM, can be naturally parallelized, we also achieve significant speed-up through massive parallelization on GPU.
@@ II-A Black-box (Data-driven) Models
The environment representation modeling from a single [23] or multiple RGB images [20] in a top-down view has been widely studied in the literature. The geometry of the scene prediction (taking into account occlusions) from stereo or depth-camera input has been addressed in [24]. The importance and benefits of sensor fusion (camera images, lidar, and radar) for environment representation have been demonstrated in [25]. The authors of RoadRunner [26] developed an end-to-end learning-based framework that predicts the traversability and elevation maps from multiple images and a lidar voxel map. However, the robot-terrain interaction and trajectory prediction tasks are not considered in these works.
@@
Figure 7: Detailed view of MonoForce’s heightmaps: Geometrical heightmap replicates lidar heightmap. Soft heightmap outlines soft parts of the terrain (those that do not generate any robot-terrain interaction forces) such as vegetation. Terrain heightmap consists of two layers: (i) Supporting terrain: geometrical heightmap with subtracted soft parts of the terrain; (ii) Friction heightmap: see significantly higher values on dirt.
@@ II-B White-box (Physics-based) Models
The vehicle-terrain interaction modeling and traction mechanics have been widely studied in the literature [33][34]. To provide accurate analysis and prediction of off-road vehicles behavior on the terrain, a set of high-fidelity physics-based software tools have been developed. For example, Vortex Studio [35] allows for realistic wheeled and tracked locomotion simulations and extensive land and even planetary environment modeling. The Chrono [36] is a physics-based engine designed for ground vehicle-tire-terrain interactions simulation. It supports the integration of various vehicle and tire models and can simulate deformable terrain. The AGX physics engine (more specifically, its terrain dynamics model agxTerrain[37] enables simulation of soil dynamics and interactions with heavy vehicles. This software delivers realistic, high-fidelity simulations for applications involving complex soil-tool interactions. However, the mentioned physics engines are not differentiable, which makes them unsuitable for end-to-end learning and integration of exteroceptive real sensor data.
@@ III Method
A detailed overview of the proposed architecture that converts images and control commands into trajectories is depicted in Figure 6. The model consists of several learnable modules that deeply interact with each other. The terrain encoder carefully transforms visual features from the input image into the heightmap space using known camera geometry. The resultant heightmap features are further refined into interpretable physical quantities that capture properties of the terrain such as its shape, friction, stiffness, and damping. Next, the physics engine combines the terrain properties with the robot model, robot state, and control commands and delivers reaction forces at points of robot-terrain contacts. It then solves the equations of motion dynamics by integrating these forces and delivers the trajectory of the robot. Since the complete computational graph of the feedforward pass is retained, the backpropagation from an arbitrary loss, constructed on top of delivered trajectories, or any other intermediate outputs is at hand.
@@ III-A Terrain Encoder
The part of the MonoForce architecture (Figure 6) that predicts terrain properties $\mathbf{m}$ from sensor measurements $\mathbf{z}$ is called terrain encoder. The proposed architecture starts by converting pixels from a 2D image plane into a heightmap with visual features. Since the camera is calibrated, there is a substantial geometrical prior that connects heightmap cells with the pixels. We incorporate the geometry through the Lift-Splat-Shoot architecture [20]. This architecture uses known camera intrinsic parameters to estimate rays corresponding to particular pixels – pixel rays, Figure 13. For each pixel ray, the convolutional network then predicts depth probabilities and visual features. Visual features are vertically projected on a virtual heightmap for all possible depths along the corresponding ray. The depth-weighted sum of visual features over each heightmap cell is computed, and the resulting multichannel array is further refined by deep convolutional network to estimate the terrain properties $\mathbf{m}$.
@@ III-D End-to-end Learning
Terrain loss that minimizes the difference between ground truth $\mathcal{H}_{t}^{\star}$ and predicted $\mathcal{H}_{t}$ supporting heightmaps containing rigid objects detected with Microsoft’s image segmentation model SEEM [32], that is derived from Segment Anything foundation model [46]:
@@
Since the architecture Figure 2 is end-to-end differentiable, we can directly learn to predict all intermediate outputs just using trajectory loss (8). An example of terrain learning with the trajectory loss is visualized in Figure 9. To make the training more efficient and the learned model explainable, we employ the geometrical loss (9) and terrain loss (10) as regularization terms. stat
@@
Figure 11: MonoForce prediction examples. Left: The robot is moving through a narrow passage between a wall and tree logs. Right: The robot is moving on a gravel road with rocks on the sides. It starts its motion from the position marked with a coordinate frame and the trajectory is predicted for $10~{}[$\sec$]$ using real control commands. The camera images are taken from the robot’s initial position (top row). The visualization includes predicted supporting terrain $\mathcal{H}_{t}$ (second row). It is additionally shown in 3D and colored with predicted friction values (third row).
@@
The Figure 11 show the prediction examples of the MonoForce model in diverse outdoor environments. From the example on the left, we can see that the model correctly predicts the robot’s trajectory and the terrain shape suppressing traversable vegetation, while the rigid obstacles (wall and tree logs) are correctly detected. The example on the right demonstrates the model’s ability to predict the robot’s trajectory ($10~{}[$\sec$]$-long horizon) with reasonable accuracy and to detect the rigid obstacles (stones) on the terrain. It could also be noticed that the surfaces that provide the robot good traction (paved and gravel roads) are marked with a higher friction value, while for the objects that might not give good contact with the robot’s tracks (walls and tree logs) the friction value is lower.
@@
We argue that the friction estimates are approximate and an interesting research direction could be comparing them with real-world measurements or with the values provided by a high-fidelity physics engine (e.g. AGX Dynamics [37]). However, one of the benefits of our differentiable approach is that the model does not require ground-truth friction values for training. The predicted heightmap’s size is $12.8\times 12.8$\mathrm{m}$^{2}$ and the grid resolution is $0.1$\mathrm{m}$$. It has an upper bound of $1~{}[$\mathrm{m}$]$ and a lower bound of $-1~{}[$\mathrm{m}$]$. This constraint was introduced based on the robot’s size and taking into account hanging objects (tree branches) that should not be considered as obstacles (Figure 17). Additionally, the terrain is predicted in the gravity-aligned frame. That is made possible thanks to the inclusion of camera intrinsics and extrinsics as input to the model, Figure 6. It also allows correctly modeling the robot-terrain interaction forces (and thus modeling the robot’s trajectory accurately) for the scenarios with non-flat terrain, for example, going uphill or downhill. This will not be possible if only camera images are used as input.
@@ IV-A Dataset: ROUGH
Figure 12: Two robot platforms used to collect the ROUGH dataset. Notice that one platform has only four massive flippers, and the other has also two main tracks.
@@
Despite the amount and annotation quality of the data provided in the RELLIS-3D sequences, it lacks examples of a robot moving over hills, obstacles, and traversing high grass. To fill this gap in autonomous off-road navigation, a new Rugged Obstacles and Uneven Ground Hardcore (ROUGH) dataset will be released, containing sensory data from forest and field scenarios recorded with the two robot platforms: MARV, Bluebotics AbsolemFigure 12. It contains several hours of driving with the mid-sized robots in challenging terrain. The ultimate goal is to provide recordings of traversals on (or through) flexible natural obstacles like tall grass, hanging tree branches, mud, dense undergrowth, etc. The ROUGH sequences are collected with shape-changing tracked robots (Figure 12), which allows capturing a much larger range of dynamic responses - simply moving the auxiliary tracks changes the center and moments of inertia and the contact surface. To correctly utilize the effects of dynamics, ROUGH provides not only sensory data but also robot models with dynamic properties. The data set also contains carefully time-synchronized point cloud scans from Ouster OS0-128 lidar and corresponding RGB images from $4$ Basler cameras installed on the robots (front, rear, left, and right). The robot localization data obtained using ICP SLAM [48] is also recorded. It was a better choice than GPS which is unreliable near or under tree canopy and close to buildings. The controls ($\mathbf{u}$ in Figure 6) in the form of commanded tracks velocities and angles are recorded and used to train and evaluate the MonoForce model.
After training the MonoForce terrain encoder on the RELLIS-3D dataset, we fine-tune it on the ROUGH data sequences. Please note that the dataset is under construction at the moment of writing and will be extended with the help of new robot platforms and data collection scenarios. The ROUGH dataset will be made publicly available [^2] https://github.com/ctu-vras/rough-dataset.